OpenStudy (anonymous):

Which point is a solution of y = 3x and x + 2y = –21? Select one: a. (–3, –9) b. (–9, –3) c. (3, –9)

OpenStudy (aripotta):

just substitute each of the coordinate in both equations to see which one works out

jimthompson5910 (jim_thompson5910):

As AriPotta is saying x + 2y = –21 x + 2(3x) = –21 ... replace y with 3x (works because y = 3x) x + 6x = -21 ... ... ... x = ???